OpenStudy (anonymous):

What are the solutions of x2 + 3x − 18 = 0? show work so i understand please?

OpenStudy (anonymous):

x^2 + 3x - 18 = 0 (x+6)(x-3) = 0 x = -6, x = 3

OpenStudy (anonymous):

x(x+3)=18 x=-6 or x=3

OpenStudy (anonymous):

x2+3x-18=0 (x+6)(x-3) - 0 x=-6, 3